Okay so I have been trying to solve this answer from my homework packet but i just dont remeber how to do it..If log(x+2)/log3-log(x-2)\log3=2, then what is x? HELP!!
Answer by stanbon(29482)

If log(x+2)/log3-log(x-2)\log3=2
------
Multiply thru by log(3) to get:
If log(x+2)-log(x-2) = 2log(3)
log[(x+2)/(x-2)] = log(9)
[(x+2)/(x-2)] = 9
Cross-multiply to solve for "x":
x+2 = 9x-18
8x = 20
x = 2.5
=================
Cheers,
Stan H.